If f: R → R is defined by f (x) = x^2 – 3x + 2, write f (f (x)). 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given,f(x)=x^2–3x+2 Then,
f(f(x))=f(x^2–3x+2)

=(x^2–3x+2)2–3(x^2–3x+2)+2,
=x^4+9x^2+4–6x^3+4x^2–12x–3x^2+9x–6+2
=x^4–6x^3+10x^2–3x

Thus,
f(f(x))=x^4–6x^3+10x^2–3x

Thanks